# Modified Axelrod Model Showing Opinion Convergence And Polarization In Realistic Scale-Free Networks

### v1

@@ IV.1 Methodology
We study the dynamics of our modified Axelrod model using Python. The scale-free network is created using the Networkx package and most of the analytical work is done using the NumPy package. Figures are produced by the Matplotlib package.
@@
Figure 2: Histograms showing the time evolution of opinion distribution with all but the parameter $t$ taken from Table 1. The $x$-axes of all subplots are the opinions in $(-1,1)$ and the $y$-axes are the occurrence frequencies of the opinion. The first six subplots have common $y$-axis limits, from 00 to $750$. Although the initial opinions concentrate around 00, the system evolves to a state of extreme opinion polarization in the long time limit in which nearly all agents have extreme opinions on all features.
We investigate the effect of each model parameter on the development of the system by varying one of them and fixing all the rest to their corresponding default values as tabulated in Table 1. We report our simulation results in histograms on the opinion distribution of feature traits. Since our initialization and evolution rules apply to different feature traits in the same way, opinion distributions of different feature traits averaged over a sufficiently large ensemble of independent runs should be the same. This is indeed what we have observed in our simulation. Consequently, to save running time, we lump the distributions of all $F$ feature traits of every agent in plotting the histogram. In each histogram, we divide the opinion into $50$ evenly spaced bins (in the usual Euclidean distance) each with a bin width of $0.04$. Each of the simulation results presented below is taken from a single run randomly selected over a sample of 10 runs. In fact, all runs give very similar distributions.
@@ IV.2 Feature Evolution
We begin by investigating the time evolution of opinions by fixing all other model parameters to the ones listed in Table 1. Figure 2 shows the time evolution of agents opinions. Clearly, opinion spreads as time progresses. When the mean number of interactions per agent $t$ is less than $\approx 80$, the opinion distribution roughly follows the Gaussian distribution whose mean is zero and the standard deviation increases gradually. When $t\gtrsim 80$, extreme opinions start to emerge. In addition, the fractional increase of extreme opinions accelerates. Eventually, extreme opinions dominate, and nearly all agents have extreme opinions on all features. That is to say, opinions become extremely polarized. This result can be qualitatively understood as follows. Since the agreement threshold $a$ is less than the standard deviation of the initial opinion $\sigma$, a sizable fraction of neighboring agents is dissimilar. Their interactions gradually lead to opinion polarization. Once a non-negligible portion of the opinion is polarized, our evolution rules accelerate further polarization for it is harder and harder to find a similar agent pair which consists of an agent with moderate opinion and another agent with an extreme opinion on the same feature. As a consequence, it is no longer possible to pull enough extreme opinion agents back to moderate ones. Thus, an extremely polarized society is formed in the end. Readers will find in Secs. IV.3IV.5 that the formation of extremely polarized opinion is a genuine and common feature of our model using reasonable parameters rather than the result of a cherry-picked set of parameters. (Surely, opinions are never polarized if $\sigma\approx 0$ or $\kappa\approx 0$. However, these sets of parameters are atypical and unrealistic.)
@@
Figure 3: Opinion distributions after $Nt/2$ time steps with $t=200$ for different values of $N$. These subplots show no significant difference in opinion distribution for different $N$.
@@
Figure 5: Opinion distribution after $Nt/2$ time steps with $N=1000$ and $t=100$ for different values of $m$. Compared with Figure 5, these subplots show that longer equilibration times are needed to reach the extremely opinion depolarization state as $m$ increases.
IV.3 Effect Of The Number Of Agents
We now study the effect of the number of agents $N$ while fixing other parameters to their default values listed in Table 1. As shown in Figure 3, there is no significant difference in the distribution between different histograms even if $N$ varies over four orders of magnitude from $10^{2}$ to $10^{5}$. This result is not surprising because, upon coarse-graining, the resultant network is still scale-free. And since the formation of extreme opinion polarization is the generic behavior of the system, the effective parameters of the renormalized system likely fall within this generic region. Finally, we find that the convergence rate of the opinion polarization is almost independent of $N$. That is to say, for a given mean number of interactions per agent $t$, the distribution between different histograms is statistically the same regardless of $N$.
IV.4 Effect Of The Average Degrees Of Agents In The Scale-free Network
@@
Here we investigate the effect of the average degree of the scale-free network $m$ on the system evolution by varying it from $4$ to $48$. As shown in Figure 4, there is no significant difference in the final steady state of the opinion distribution among all agents. They all evolve to a highly polarized opinion state.
@@
To study the rate of convergence to this highly polarized state, we reduce the average number of interactions per agent $t$ from $200$ to $100$. Figure 5 depicts that the equilibration time to the extremely depolarized state increases as $m$ increases. This result can be understood as follows. An agent has more neighbors on average as $m$ increases. Consequently, the opinion of an agent is influenced by more neighbors each has statistically independent opinions on each feature trait initially. Thus, it takes a longer time for opinions of the system to polarize.
Figure 6: Opinion distribution after $Nt/2$ time steps for $N=1000$ and $t=200$ for different values of $a$. Subplots (a)–(h) depict that opinions of the equilibrated states are extremely polarized when $a\lesssim\sigma$. Subplots (j)–(l) show that the system reaches consensus of a single moderate opinion in the long run if $a\gg\sigma$. In between these two parameter ranges, both moderate and polarized opinions coexist in the long run as in subplot (i).
IV.5 Effect Of The Agreement Threshold
@@
We now study the effect of agreement threshold $a$ on the system. When $a\lesssim\sigma$, as shown in subplots (a)–(h) in Figure 6, the system equilibrates to two dominating peaks at a polarized opinion as in the generic case. This is because the initial opinion distribution is sufficiently diverse to allow a sizable portion of neighboring dissimilar agents. Their interactions seed more extreme opinions. In contrast, when $a\gg\sigma$, the system evolves to a single peak of moderate opinion and the extreme opinions population is very small. (See subplots (j)–(l) in Figure 6.) The reason is that with a large value of $a$, almost all agent pairs are similar. This leads to opinion convergence for at least a majority portion of the population. For an intermediate value of $a$, as depicted in Figure 6(i), the coexistence of a non-negligible number of extreme and moderate opinions is observed.
@@
In summary, we find that the value of $a/\sigma$ is the most important factor determining the asymptotic behavior of the system. It is natural to expect that $a<\sigma$ in human societies for otherwise the boomerang effect can hardly be observed. The implication of our modified Axelrod model is that societies will evolve toward extreme opinion polarization — a phenomenon in our contemporary society that leads to all kinds of troubles making most people unhappy and some to go violence.
Figure 7: Evolution of opinion distribution for different portion of sympathetic agents $p_{\text{s}}$ and mean number of interactions per agent $t$. All other parameters are set to default values listed in Table 1.
Figure 8: Evolution of opinion distribution for different values of $p_{\text{s}}$ after the system has reached a steady final state. All parameters taken are the same as those in Figure 7.
IV.6 Effects Of Other Parameters
@@
Last but not least, we consider the effects of varying each of the following parameters — $F$, $\kappa$ and $\mu$. (There is no need to consider the variations of $\bar{x}$ and $\sigma$. This is because the opinion evolution rule in Step 3 of our model depends only on the hyperbolic distances between opinion pairs. Changing the value of $\bar{x}$ simply shifts the opinion distribution by $\bar{x}$ in the hyperbolic distance metric. Besides, rescaling $\sigma\mapsto s\sigma$ and $a\mapsto sa$ just multiply the hyperbolic distance between the resultant $X_{f}$’s by the same factor $s$. Thus Step 3 implies that the resultant opinion distributions are scaled by the factor $s$ in hyperbolic distance. These two changes do not affect the number of agents with moderate as well as extreme opinions.) We find that varying $\kappa$ alters the final opinion distribution while varying the others do not. This is consistent with the analysis above. The reason is that by varying $\kappa$, we increase the fraction of similar agents and hence slow down the emergence of polarized opinion. Therefore, the simulation result is very similar to the result shown in Figure 6. We do not show the corresponding histograms here to save space.
V Reducing Opinion Polarization By Introducing Sympathetic Agents
@@
Could we make the opinion distribution of the eventual states less polarized within the framework of our modified Axelrod model? We would like to investigate possible interventions that could alleviate extreme opinion polarization. As a general principle, we would like to make our interventions as small as possible, hopefully in the form of making small changes in our rules. This is because changing human psychological behavior en masse is not easy, at least in the short term. Only these kinds of small changes have a real chance to occur in practice. Here we propose to alter the behavior of a small fraction of agents. Specifically, we fix a small portion $p_{\text{s}}$ of the agents to be sympathetic. A sympathetic agent always interacts with its neighbors using the rules of a similar agent. One way to think of it is that each agent has a different agreement threshold, and that of a sympathetic agent is arbitrarily large. Their readiness to be influenced and persuaded by all kinds of opinions is the key to reaching a consensus in the population.

### v2

@@ IV.1 Methodology
We study the dynamics of our modified Axelrod model using Python [52]. The scale-free network is created using the Networkx package and most of the analytical work is done using the NumPy package. Figures are produced by the Matplotlib package.
@@
Figure 1: Histograms showing the time evolution of opinion distribution with all but the parameter $t$ taken from Table 1. The $x$-axes of all subplots are the opinions in [1,1][-1,1] and the $y$-axes are the occurrence frequencies of the opinion. Although the initial opinions concentrate around 0, the system evolves to a state of extreme opinion polarization in the long time limit in which nearly all agents have extreme opinions on all features. Note that the $y$-axis ranges of all histogram subplots in this paper may be different.
We investigate the effect of each model parameter on the development of the system by varying one of them and fixing all the rest to their corresponding default values as tabulated in Table 1. Each of the simulation results presented below is taken from a single run randomly selected over a sample of 10 runs. And a new scale-free network is used for each run. In fact, all runs give very similar distributions. We report our simulation results in histograms on the opinion distribution of feature traits. Since our initialization and evolution rules apply to different feature traits in the same way, opinion distributions of different feature traits averaged over a sufficiently large ensemble of independent runs should be the same. This is indeed what we have observed in our simulation. Consequently, to save running time, we lump the distributions of all $F$ feature traits of every agent in plotting the histogram. In each histogram, we divide the opinion into $50$ evenly spaced bins (in the usual Euclidean distance) each with a bin width of $0.04$.
@@ IV.2 Feature Evolution
We begin by investigating the time evolution of opinions by fixing all other model parameters to the ones listed in Table 1. Fig. 1 shows the time evolution of agents opinions. Clearly, opinion spreads as time progresses. When the mean number of interactions per agent $t$ is less than $\approx 20$, the opinion distribution roughly follows the Gaussian distribution whose mean is zero and the standard deviation increases gradually. When $t\gtrsim 20$, extreme opinions start to emerge. In addition, the fractional increase of extreme opinions accelerates. Eventually, extreme opinions dominate, and nearly all agents have extreme opinions on all features. That is to say, opinions become extremely polarized.
@@
Figure 2: Opinion distributions after $Nt/2$ time steps with $t=200$ for different values of $N$. These subplots show no significant difference in opinion distribution for different NN.
Our simulation result can be qualitatively understood as follows. Since the agreement threshold $a$ is less than the standard deviation of the initial opinion $\sigma$, a sizable fraction of neighboring agents is dissimilar. Their interactions gradually lead to opinion polarization. Once a non-negligible portion of the opinion is polarized, our evolution rules accelerate further polarization for it is harder and harder to find a similar agent pair which consists of an agent with moderate opinion and another agent with an extreme opinion on the same feature. As a consequence, it is no longer possible to pull enough extreme opinion agents back to moderate ones. Thus, an extremely polarized society is formed in the end. Readers will find in our subsequent analysis that the formation of extremely polarized opinion is a genuine and common feature of our model using reasonable parameters rather than the result of a cherry-picked set of parameters. (Surely, opinions are never polarized if $\sigma\approx 0$ or $\kappa\approx 0$. However, these sets of parameters are atypical and unrealistic.)
Let us compare our opinion distribution results with several models on market. The three closest models on market are the ones reported by Chau et al. [13], Jager and Amblard [27] as well as Liu et al. [36]. They used Euclidean metric bounded confidence with assimilation as well as contrast dynamics. Agents are placed on a complete graph in Refs. [13, 27], regular 2-d lattice in Ref. [27] and a ring in Ref. [36]. Using suitable parameters, the eventual agent opinion distributions in these two models exhibit bipolar extreme polarization very similar to ours.
IV.3 Effect Of The Number Of Agents
We now study the effect of the number of agents $N$ while fixing other parameters to their default values listed in Table 1. As shown in Fig. 2, there is no significant difference in the distribution between different histograms even if $N$ varies over four orders of magnitude from $10^{2}$ to $10^{5}$. This result is not surprising because, upon coarse-graining, the resultant network is still scale-free. And since the formation of extreme opinion polarization is the generic behavior of the system, the effective parameters of the renormalized system likely fall within this generic region. Finally, we find that the convergence rate of the opinion polarization is almost independent of $N$. That is to say, for a given mean number of interactions per agent $t$, the distribution between different histograms is statistically the same regardless of $N$.
Figure 3: Opinion cluster distributions of a typical run of our model on (a) a B-A network and (b) $20\times 20$ lattice using parameters in Table 1 except that $N=400$. Agents in the same opinion cluster are represented by dots of the same color; and no adjacent cluster is of the same color. Each adjacent agent pair in (a) is linked by a solid black line. In (b), a thick solid black (thin dashed black) line is drawn between a pair of adjacent similar (dissimilar) agents.
IV.4 Distributions Of Opinion Clusters And Their Sizes
We turn to study the opinion cluster distribution and its relation with $N$ where an opinion cluster is defined to be a maximal (in the sense of set inclusion) path-connected set of agents such that any pair of adjacent agents (with respect to the underlying network topology) in the set is similar. Fig. 3(a) shows the distribution of opinion clusters for a typical run using parameters in Table 1 at $t=200$ with $N$ replaced by $400$ to avoid over-crowding the graph. We find that in about $2/3$ of the runs, one dominant opinion cluster consisting of about 50% to 60% of the population shows up. (One example is Fig. 3(a) whose dominant opinion cluster is made up of the pale yellow dots in the central region of the figure.) Whereas for the remaining $1/3$ of the runs, one large and one much smaller dominant opinion clusters appear along side with many small clusters. Besides, Fig. 4(a) tells us that as $N$ increases, both the chance of having more than one dominant opinion cluster and the standard deviation of the size of the dominant opinion cluster decrease. For instance, when $N=10^{5}$, the distribution of the size of the dominant opinion cluster sharply peaks in the interval $[0.54N,0.56N]$ up to $3\sigma$ level. Furthermore, for all the number of agents $N$ used in our simulation, every remaining cluster contains much fewer number of agents. In other words, there is no intermediate sized opinion cluster in the system. To summarize, a highly fragmented society with a significant number of highly connected agents holding similar extreme opinions is formed. More importantly, numerous small clusters many of them holding opposite extreme opinions are also present. This polarization pattern resembles those observed in the real world such as the great opinion differences between those living in cities and in rural areas as well as the emergence of political extremism due to distrust of the elites. It is instructive to explore the potential of applying our model to study this type of society as the general trend reported above is generic over typical sets of parameters. In contrast, Fig. 3(b) shows the opinion clusters of a typical run of our model when agents are placed in a $20\times 20$ lattice using the same parameters as in Fig. 3(a). Unlike the case of using B-A network, here we see a much greater number of similarly sized opinion clusters. Besides, no dominant cluster is present. (For instance, the largest opinion cluster in Fig. 3(b) consists of $4$ agents represented by purple dots near the bottom right corner of the figure.)
Figure 4: The average number of opinion cluster divided by the number of agents $n/N$ as a function of cluster size $s$ and number of agents $N$ on a B-A network in (a) log-log and (b) semi-log plots as well as on a square lattice in (c) log-log and (d) semi-log plots using parameters in Table 1 except that $N$ can vary. The red dots, green squares and blue triangles are for $N=10^{3}$, $10^{4}$ and $10^{5}$, respectively for the B-A network and the responding closest square number for the square lattice. Each set of data is obtained by averaging over $1024$ independent runs. The uncertainty of $\log_{10}[n(s,N)/N]$ of each data point is of order of $(NS)^{-1/2}/n$ and the corresponding error bar is omitted in the subfigures for clarity. The dashed line in each subfigure is the best fit curve whose parameters are shown in Table 2.
We further investigate the opinion cluster size distribution. Let us denote the average number of opinion cluster of size $s$ in a system of $N$ agents by $n(s,N)$. Figs. 4(a) and (b) show the log-log and semi-log plots of $n(s,N)/N$. The data points are obtained for $N=10^{3},10^{4}$ and $10^{5}$ each averaged over at least $S=1024$ runs using parameters listed in Table 1 except that $N$ can vary. Scaling is observed in $n(s,N)/N$ as curves for different values of $N$ collapse. Moreover,
$\frac{n(s,N)}{N}\approx\cases{A}s^{-\tau}&\text{for~}s\lesssim 10,\\ Be^{-\lambda s}&\text{for~}15\lesssim s\lesssim 35,$ (7)
with fitting parameters $A,\tau,B$ and $\lambda$ tabulated in Table 2. Figs. 4(a) and (b) depict that opinion cluster distribution follows a power law over roughly one decade for $s\lesssim 10$ and an exponential decay law over roughly two decades for $15\lesssim s\lesssim 35$. The power law distribution for small cluster size can be understood as follows. Since $n(s,N)$ obeys the sum rule $\sum_{s}s\ n(s,N)=N$ and agents are placed on a scale-free network, we expect that $n(s,N)/N\sim s^{-\tau}$ for sufficiently small $s$.
Parameter B-A network square lattice
$\log_{10}A$ $51\pm 0.02 0.0$$02\pm 0.06$
$\tau$ $3.2\pm 0.2$ $5.9\pm 0.4$
$\log_{10}B$ $2.4\pm 0.1$ $2.9\pm 0.1$
$\lambda$ $0.16\pm 0.01$ $0.19\pm 0.01$
Table 2: Best fit parameters for the subplots in Fig. 4.
Next, we explore the effect of the underlying network topology on opinion cluster size distribution. Figs. 4(c) and (d) show the log-log and semi-log plots of $n/N$ as a function of $s$ for different $N$ when agents are placed on a square lattice. Just like the B-A network case, scaling is observed as the curves collapse for any $N$. But unlike the B-A network situation, power law is observed in a much smaller region of $s\lesssim 4$ — a region too small to be of statistical significance. Furthermore, exponential decay is found for $5\lesssim s\lesssim 25$. The corresponding best fit parameters are listed in Table 2. Therefore, it appears that the use of scale-free network enlarges the power law regime of opinion cluster distribution.
We now compare our opinion cluster size distribution results with a few existing models in the literature. Opinion cluster size distribution of our model in B-A network is similar to that by Jacobmeier [29] that also used B-A network but only with assimilative dynamics. Both shows power law distribution but with very different exponents over very different regions — their exponent is close to 0 and is valid for intermediate values of $s$ over a range of about $1.5$ decades while our exponent is about $3.4$ and is valid for $s\lesssim 10$. In contrast, their model shows clustering in opinions. Furthermore, our model using B-A network shares similar opinion cluster size distribution with the model by Deffuant et al. [7] that used assimilative dynamics only in a regular 2-d lattice. However, their distribution exhibits power law but with no obvious scaling. Our 2-d square lattice result is also similar to the stable configuration of the polarization phase of the J-A model [27]. Besides, ours exhibits scaling but without any power law. In addition, the behavior of our model is different from that in Dinkelberg et al. [30] as the cluster size distribution of the later is weakly dependent on the underlying network used.
@@ IV.4 Distributions Of Opinion Clusters And Their Sizes
IV.5 Effect Of The Average Degrees Of Agents In The Scale-free Network
@@
Here we investigate the effect of the average degree of the scale-free network $m$ on the system evolution by varying it from $4$ to $48$. As shown in Fig. 5, there is no significant difference in the final steady state of the opinion distribution among all agents. They all evolve to a highly polarized opinion state.
Figure 6: Opinion distribution after $Nt/2$ time steps with $N=1000$ and $t=100$ for different values of $m$. Compared with Fig. 5, these subplots show that longer equilibration times are needed to reach the extremely opinion depolarization state as $m$ increases.
@@
To study the rate of convergence to this highly polarized state, we reduce the average number of interactions per agent $t$ from $200$ to $100$. Fig. 6 depicts that the equilibration time to the extremely depolarized state increases as $m$ increases. This result can be understood as follows. An agent has more neighbors on average as $m$ increases. Consequently, the opinion of an agent is influenced by more neighbors each has statistically independent opinions on each feature trait initially. Thus, it takes a longer time for opinions of the system to polarize.
IV.6 Effect Of The Agreement Threshold
@@
We now study the effect of agreement threshold $a$ on the system. When $a\lesssim\sigma$, as shown in subplots (a)–(e) in Fig. 7, the system equilibrates to two dominating peaks at a polarized opinion as in the generic case. This is because the initial opinion distribution is sufficiently diverse to allow a sizable portion of neighboring dissimilar agents. Their interactions seed more extreme opinions. In contrast, when $a\gtrsim\sigma$, the system evolves to a single peak of moderate opinion and the extreme opinions population is very small. (See subplots (h)–(l) in Fig. 7.) The reason is that with a large value of $a$, almost all agent pairs are similar. This leads to opinion convergence for at least a majority portion of the population. For $a\approx\sigma$, as depicted in subplots (f) and (g) in Fig. 7, the coexistence of a non-negligible number of extreme and moderate opinions is observed.
Figure 7: Opinion distribution after $Nt/2$ time steps for $N=1000$ and $t=200$ for different values of $a$. It shows that opinion polarization occurs only when $a\lesssim\sigma$.
@@
In summary, we find that the value of $a/\sigma$ is the most important factor determining the asymptotic behavior of the system. It is natural to expect that $a<\sigma$ in human societies for otherwise the repulsive dynamics can hardly be observed. The implication of our modified Axelrod model is that societies will evolve toward extreme opinion polarization — a phenomenon in our contemporary society that leads to all kinds of troubles making most people unhappy and some to go violence.
IV.7 Effects Of Other Parameters
@@
Last but not least, we consider the effects of varying each of the following parameters — $F$, $\kappa$ and $\mu$. We find that varying $\kappa$ alters the final opinion distribution while varying the others do not. This is consistent with the analysis above. The reason is that by varying $\kappa$, we increase the fraction of similar agents and hence slow down the emergence of polarized opinion. Therefore, the simulation result is very similar to the result shown in Fig. 7. We do not show the corresponding histograms here to save space.
Surely, we may further consider the effects of $\bar{x}$ and $\sigma$. However, we are not going to do so here. Setting a non-zero $\bar{x}$ introduces initial bias on opinion distribution that may not be easy to justify. Whereas if $\sigma\ll 1$, by rescaling the $X_{f}$, it is clear that the ratio $a/\sigma$ is a determining factor of the system dynamics. As we have already investigated the effect of $a$ on the dynamics, this makes the study of $\sigma$ redundant.
V Reducing Opinion Polarization By Introducing Empathetic Agents
@@
Could we make the opinion distribution of the eventual states less polarized within the framework of our modified Axelrod model? We would like to investigate possible interventions that could alleviate extreme opinion polarization. As a general principle, we would like to make our interventions as small as possible, hopefully in the form of making small changes in our rules. This is because changing human psychological behavior en masse is not easy, at least in the short term. Only these kinds of small changes have a real chance to occur in practice. Here we propose to alter the behavior of a small fraction of agents. Specifically, we fix a small portion $p_{\text{s}}$ of the agents randomly chosen among those whose number of neighbors in the B-A network rank top $100\rho_{\text{s}}$% to be empathetic. A empathetic agent always interacts with its neighbors using the rules of a similar agent. One way to think of it is that each agent has a different agreement threshold, and that of a empathetic agent is arbitrarily large. Their readiness to be influenced and persuaded by all kinds of opinions is the key to reaching a consensus in the population.
Figure 8: Evolution of opinion distribution for different portion of empathetic agents $p_{\text{s}}$ and mean number of interactions per agent $t$ when $\rho_{\text{s}}=1$. All other parameters are set to default values listed in Table 1.
Figure 9: Evolution of opinion distribution for different values of $p_{\text{s}}$ after the system has almost reached a steady final state. All parameters taken are the same as those in Fig. 8.
Figure 10: Evolution of opinion distribution for different portion of empathetic agents $p_{\text{s}}$ and mean number of interactions per agent $t$ when $\rho_{\text{s}}=0.3$. All other parameters are set to default values listed in Table 1.
Figure 11: Evolution of opinion distribution for different values of $p_{\text{s}}$ after the system has almost reached a steady final state. All parameters taken are the same as those in Fig. 10.
V.1 The Case Of Unbiased Empathetic Agent Selection